### v1

# Reinforced Diffuser for Red Teaming Large Vision-Language Models

Abstract
The rapid advancement of large Vision-Language Models (VLMs) has raised significant safety concerns, particularly regarding their vulnerability to jailbreak attacks. While existing research primarily focuses on VLMs’ susceptibility to harmful instructions, this work identifies a critical yet overlooked vulnerability: current alignment mechanisms often fail to address the risks posed by toxic text continuation tasks. To investigate this issue, we propose a novel Red Team Diffuser (RTD) framework, which leverages reinforcement learning to generate red team images that effectively induce highly toxic continuations from target black-box VLMs. The RTD pipeline begins with a greedy search for high-quality image prompts that maximize the toxicity of VLM-generated sentence continuations, guided by a Large Language Model (LLM). These prompts are then used as input for the reinforcement fine-tuning of a diffusion model, which employs toxicity and alignment rewards to further amplify harmful outputs. Experimental results demonstrate the effectiveness of RTD, increasing the toxicity rate of LLaVA outputs by 10.69% on the original attack set and 8.91% on a hold-out set. Moreover, RTD exhibits strong cross-model transferability, raising the toxicity rate by 5.1% on Gemini and 26.83% on LLaMA. These findings reveal significant deficiencies in existing alignment strategies, particularly their inability to prevent harmful continuations. Our work underscores the urgent need for more robust and adaptive alignment mechanisms to ensure the safe deployment of VLMs in real-world applications.
Disclaimer: This paper contains content that may be disturbing or offensive.
footnotetext: $\dagger$  Corresponding author.
1 Introduction
The widespread deployment of large Vision-Language Models (VLMs), such as OpenAI’s GPT-4o Achiam et al. (2023) and Google’s Gemini Team et al. (2023), has raised significant concerns about their safety to potential jailbreak attacks. While VLMs inherit alignment properties from Large Language Models (LLMs), their multimodal nature introduces novel safety risks. Current jailbreak attacks on VLMs leverage either unimodal or multimodal approaches to manipulate the models into responding to harmful requests, such as providing bomb-making instructions Wang et al. (2024a).
However, research remains limited on the toxicity of generated outputs, such as insults and offensive language. For responsible deployment, it is crucial to rigorously safeguard these models against producing any harmful or unsafe content.
In this work, we demonstrate that existing safety mechanisms fail to prevent the generation of offensive, discriminatory, or hateful language when VLMs encounter biased or misleading contexts in text continuation tasks. In such scenarios, they often produce continuations containing harmful expressions, exposing a critical vulnerability.
Figure 1: Example of a successful jailbreak on LLaVA. The proposed Red Team Diffuser (RTD) generates toxic-triggering images despite using the same prompt as Stable Diffusion.
As multimodal models, the generated content of VLMs can be significantly influenced by visual modalities, despite having the same text input. Identifying which images can effectively serve as triggers for toxic continuations remains an open challenge. Building on the substantial progress made in white-box jailbreaks Qi et al. (2024); Wang et al. (2024a), one potential approach for triggering toxic continuation is the use of adversarial images. However, optimization-based adversarial images require full access to VLM parameters, which is impractical in real-world applications, especially in VLM-as-a-Service scenarios where the VLM operates as a black-box system. Additionally, such adversarial images often exhibit noticeable distortions, making them susceptible to detection and defense mechanisms Nie et al. (2022); Zhang et al. (2023); Zhao et al. (2024). Moreover, studies have demonstrated that adversarial images typically show limited transferability between VLMs, even when crafted to target multiple VLMs simultaneously Schaeffer et al. .
These limitations motivate us to explore alternative black-box attack strategies that do not rely on full model access and generate less detectable perturbations. Current black-box methods typically rely on manually crafted pipelines to generate image-text pairs, often involving typographic attacks and the use of pre-trained diffusion models to generate query-relevant images Gong et al. (2023); Liu et al. (2023); Teng et al. (2024). However, these approaches suffer from several drawbacks. First, structure-based attacks result in a restricted attack surface, producing predictable patterns that can be easily mitigated by defense mechanisms Wang et al. (2024c). Second, such images are primarily designed to bypass sensitive word detection and are not well-suited for triggering harmful continuations in our task. Recently, automated red-teaming approaches for VLMs have gained attention. IDEATOR Wang et al. (2024b) introduced the concept of “VLMs against VLMs”, employing a red team VLM to generate jailbreak text prompts paired with images created using pre-trained diffusion models Rombach et al. (2022). While this represents significant progress, it raises critical questions about the effectiveness of pre-trained diffusion models in jailbreaking tasks. Are these pre-trained models truly optimal for inducing harmful content in VLMs? Can images generated directly by diffusers reliably trigger such behavior? To answer these questions, we explore the potential of reinforced diffusers, which leverage reinforcement fine-tuning to optimize red team image generation for more effective and targeted VLM jailbreaking.
In this paper, we propose Red Team Diffuser (RTD), a novel diffusion model framework designed to generate unrestricted images that effectively bypass safety mechanisms in VLMs. Our approach leverages a two-stage process to systematically exploit vulnerabilities in VLMs. First, we utilize an LLM to perform a greedy search for image prompts that maximize the toxicity of VLM responses, identifying highly effective image prompts for red teaming purposes. Second, we fine-tune a diffusion model through reinforcement learning, jointly optimizing a toxicity reward to amplify the harmfulness of generated outputs and an image-prompt alignment reward to ensure consistency between prompts and images as a regularization term. Figure 1 illustrates the effectiveness of our proposed method in attacking LLaVA Liu et al. (2024b), showcasing its ability to systematically induce harmful outputs. This framework reveals critical vulnerabilities in VLMs, underscoring the need for further research to address these safety challenges.
The main contributions of our work can be summarized as:
  • We introduce Red Team Diffuser (RTD), a novel framework that, to the best of our knowledge, is the first to investigate the use of unrestricted images to effectively induce harmful outputs from VLMs.
  • We propose a novel two-stage black-box attack strategy, First, an LLM performs a greedy search to identify image prompts that maximize the toxicity of VLM responses. Second, a diffusion model is fine-tuned through reinforcement learning, leveraging toxicity and alignment rewards to enhance harmful outputs.
  • Experiments demonstrate that RTD increases LLaVA’s toxicity rate by 10.69% on the original attack set. Furthermore, RTD exhibits strong transferability across prompts and models, achieving an 8.91% increase on a hold-out set, as well as toxicity rate gains of 5.1% on Google’s Gemini and 26.83% on Meta’s Llama.
2 Related Work
2.1 Large Vision-Language Models
Large Vision-Language Models (VLMs) extend the functionality of Large Language Models (LLMs) by integrating visual modalities via an image encoder and modality alignment mechanisms. For example, MiniGPT-4 Zhu et al. (2023) aligned a frozen vision encoder Dosovitskiy et al. (2020) with a frozen Vicuna Chiang et al. (2023) using a single linear projection layer. InstructBLIP Dai et al. (2023), based on BLIP-2 Li et al. (2023b), introduced an instruction-tuning paradigm that leverages additional instructions to extract more useful visual features. LLaVA Liu et al. (2024b) used the text-only GPT-4 Achiam et al. (2023) to generate multimodal text-image instruction-following data and served as an end-to-end trained model that connects a vision encoder Radford et al. (2021) with LLaMA Touvron et al. (2023). While these models demonstrate significant potential in multimodal tasks, the integration of an additional visual modality introduces new avenues for attacks, distinct from those in text-only LLMs Shayegani et al. (2023b); Liu et al. (2024a); Li et al. (2023a); Ma et al. (2025).
Figure 2: Overview of the proposed method. Given an incomplete sentence, Gemini is used to generate an image prompt designed to maximize the likelihood of toxic continuation. The diffusion model generates an image based on this prompt. Toxicity and alignment rewards are then computed from the target VLM (LLaVA) response and used to reinforce fine-tuning of the diffusion model.
2.2 Jailbreak Attacks against VLMs
Jailbreak attacks against VLMs exploit the interaction between visual and textual inputs to bypass safeguards, enabling the generation of harmful content. These attacks can be categorized as either white-box or black-box methods. White-box jailbreak attacks Bagdasaryan et al. (2023); Bailey et al. (2023); Shayegani et al. (2023a); Carlini et al. (2024); Qi et al. (2024); Niu et al. (2024); Wang et al. (2024a) typically leverage gradient information to perturb input images or text to increase the likelihood of harmful outputs. While effective, they require full or partial access to model parameters, making them less practical and more easily detectable Nie et al. (2022); Zhang et al. (2023). In contrast, black-box jailbreak attacks present more practical and scalable solutions, as they do not require direct access to model parameters but instead exploit external vulnerabilities. Despite their advantages, many existing black-box approaches rely heavily on typography-based manipulations, making them susceptible to detection techniques like OCR. For example, FigStep Gong et al. (2023) converts harmful text into images using typography to bypass safety mechanisms. Similarly, approaches like MM-SafetyBench and VRP Liu et al. (2023); Ma et al. (2024) also employ typography-driven attacks as part of their methods, compromising their stealthiness against advanced detection systems. To address these limitations, the concept of red team VLMs has advanced with IDEATOR Wang et al. (2024b), which autonomously generates jailbreak image-text pairs in a black-box manner. By providing a scalable and efficient solution for creating adversarial inputs, IDEATOR has demonstrated the potential of red teaming to expose vulnerabilities in VLMs. However, while IDEATOR focuses on leveraging red team VLMs, our work broadens the scope of red teaming by exploring the feasibility of red team diffusers, which are orthogonal to red team VLMs and pose greater challenges. Specifically, we investigate the capability of unrestricted images to trigger harmful content generation, highlighting the significant threats posed by such inputs to VLMs. Compared to existing approaches, our work presents a more challenging scenario by addressing the limitations of typography-based attacks and provides a rigorous framework for evaluating the robustness of VLMs against adversarial inputs.
3 Preliminary
Denoising Diffusion Policy Optimization  RTD employs the Denoising Diffusion Policy Optimization (DDPO) algorithm Black et al. (2023) to fine-tune a pretrained diffusion model Rombach et al. (2022). Unlike traditional supervised learning with labeled data, DDPO leverages reinforcement learning (RL) to optimize the model by framing the iterative denoising process as a multi-step Markov Decision Process (MDP). At each step, actions (denoising predictions) are chosen to maximize the cumulative reward. The state of the MDP at time step $t$ is defined as $\mathbf{s}_{t}=\left(\mathbf{c},t,\mathbf{x}_{T-t}\right)$, where $\mathbf{c}$ represents the context, $t$ is the current time step, $\mathbf{x}_{T-t}$ is the noisy sample at time step $t$, and $T$ is the total number of denoising steps. Note that the MDP is a denoising process, i.e., the reverse diffusion process that recreates the true sample $\mathbf{x}_{0}$ from a Gaussian noise input $\mathbf{x}_{T}\sim\mathcal{N}(0,1)$. The agent (i.e., the diffuser)’s action at time step $t$ is defined as $a_{t}=\mathbf{x}_{T-t-1}$, which corresponds to the predicted previous state (i.e., the denoised sample) in the reverse process. The agent’s policy is thus $\pi_{\theta}\left(\mathbf{a}_{t}\mid\mathbf{s}_{t}\right)=p_{\theta}\left(\mathbf{x}_{T-t-1}\mid\mathbf{x}_{T-t},\mathbf{c}\right)$. The task reward for the agent $R(\mathbf{s}_{t},\mathbf{a}_{t})$ is structured as $R(\mathbf{s}_{t},\mathbf{a}_{t})=r(\mathbf{x}_{0},\mathbf{c})\text{ if }t=T\text{ else }0$. That is, the reward is non-zero only at the final time step ($t=T$), and the cumulative reward of a trajectory is equal to $r(\mathbf{x}_{0},\mathbf{c})$.
The goal of DDPO is to maximize the expected cumulative reward, which is defined over the sample distribution $p_{\theta}(x_{0}\mid c)$ and the context distribution $p(c)$:
$\mathcal{J}(\theta)=\mathbb{E}_{\mathbf{c}\sim p(\mathbf{c}),\mathbf{x}_{T}\sim\mathcal{N}(0,1),a\sim\pi_{\theta}(\cdot|s)}\left[R\left(s,a\right)\right].$
This objective encourages the model to learn a policy that generates samples close to the target distribution. To optimize the objective across multiple steps, we employ an importance sampling-based policy gradient estimator Kakade and Langford (2002):
$\nabla_{\theta}\mathcal{J}=\mathbb{E}\Bigg{[}\sum_{t=0}^{T}\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})}\nabla_{\theta}\log\pi_{\theta}(a_{t}|s_{t})R\left(s,a\right)\Bigg{]},$
where the expectation is taken over denoising trajectories generated by the previous policy $\pi_{\theta_{\text{old}}}$. This estimator leverages the ratio of the current model’s likelihood to the previous model’s likelihood, allowing for more efficient updates.
4 Methodology
4.1 Threat Model
Attack Goals  We focus on the sentence continuation task, an underexplored area in alignment mechanisms that poses unique challenges due to the open-ended nature of model outputs. The attacker aims to bypass the model’s safety measures and induce toxic continuations, such as the generation of obscene or profane language. Toxic continuations can also be viewed as a form of multimodal prompt injection, where carefully designed text prefixes and image triggers manipulate the model’s output into generating harmful content.
Adversary Capabilities  We assume the attacker has only black-box access to the victim VLM, simulating real-world interactions via external interfaces. By reinforcement fine-tuning red team diffusers based on the VLM’s feedback, RTD learns the VLM’s behavior patterns and identifies vulnerabilities, refining its policy to trigger toxic content.
4.2 Greedy Prompt Search
A central challenge in our task is the design of effective image prompts to trigger harmful content. To address this challenge and integrate the toxic continuation task into a multimodal framework, we utilize Google’s Gemini Team et al. (2023), disabling its safety settings to generate image prompts for each given sentence. Specifically, we instruct Gemini to produce image prompts that preserve the original sentence’s semantics while amplifying harmful themes, thereby increasing the likelihood of generating toxic content. Additionally, we leverage in-context learning Brown (2020) by providing examples of previously identified harmful sentence continuations to guide Gemini in generating image prompts consistent with these harmful patterns.
However, a key limitation arises from the fact that image prompts generated by LLMs may not be optimal for eliciting highly toxic continuations. To address this, we propose a greedy prompt search algorithm designed to identify near-optimal image prompts. For each sentence, the algorithm iteratively generates image prompts using Gemini and evaluates their associated toxicity scores. The toxicity score of each new prompt is compared with the previous one: if the new score is higher, the iteration continues; otherwise, the search terminates.
4.3 Reward Functions for RTD
We now introduce the dual-reward mechanism for RTD, which consists of a toxicity reward and an alignment reward.
Toxicity Reward: In the toxic continuation task, we prompt the target VLM $\mathcal{M}$ to extend a given sentence $\mathbf{y}$, which is paired with an image $\mathbf{x}$ generated by the RTD. The toxicity of the generated continuation is then assessed using the Detoxify classifier Hanu and Unitary team (2020), which evaluates toxicity across six attributes, with scores ranging from 0 (least toxic) to 1 (most toxic). The average of these scores is used as the reward function to guide the diffusion model’s optimization process:
$\mathcal{R}_{\text{tox}}(\mathcal{M}(\mathbf{x},\mathbf{y}))=\frac{1}{6}\sum_{i=1}^{6}\text{toxicity}_{i}(\mathcal{M}(\mathbf{x},\text{``Please}\\ \text{``extend the following sentence: "}+\mathbf{y}))$
Alignment Reward: Following Black et al. (2023), we introduce an alignment reward based on BERTScore Zhang et al. (2019) to improve multimodal understanding. In the context of our black-box attack, we expect the generated image $\mathbf{x}$ to be correctly interpreted by the target VLM and aligned with the intended semantics of the given image prompt $p$. Furthermore, we treat this alignment reward as a regularization term, helping to mitigate the risk of model collapse. Specifically, we define the alignment reward as:
$\mathcal{R}_{\text{align}}(p,\mathcal{M}(\mathbf{x}))=\frac{1}{|p|}\sum_{i=1}^{|p|}\max_{j}\big{(}S(p_{i},\mathcal{M}(\mathbf{x},\\ \text{``Describe the image in short."})_{j})\big{)}$
[…]
Total Reward: The total reward is a weighted sum of the toxicity reward and the alignment reward, controlled by a hyperparameter $\lambda$:
$\mathcal{R}_{\text{total}}=\mathcal{R}_{\text{tox}}+\lambda\mathcal{R}_{\text{align}}$
This combined reward allows the RTD to generate both toxic and semantically aligned image, thus achieving a balance between maximizing the harmfulness and maintaining image coherence. The overview of RTD is illustrated in Figure 2, while the detailed strategy is summarized in Algorithm 1.
Algorithm 1 Training Procedure of Red Team Diffuser
Require: Pre-trained Diffuser $D_{\theta}$, frozen LLM (e.g., Gemini) $\mathcal{G}$, target VLM $\mathcal{M}$, sentence continuation corpus $Y:=\{\mathbf{y}^{i}\}_{i=1}^{m}$, batch size $b$, learning rate $\alpha$, policy update steps $N$;
1:  Initialize $D_{\theta}$ with pre-trained weights and $P=\{\}$.
2:  for each sentence $\mathbf{y}^{i}\in Y$ do
3:     Initialize toxicity score $\mathcal{T}^{\text{max}}=0$.
4:     Query $\mathcal{G}$ with $\mathbf{y}^{i}$ to generate initial image prompt $p^{i}$.
5:     Generate initial image $\mathbf{x}^{i}=D_{\theta}(p^{i})$.
6:     Query $\mathcal{M}$ with $(\mathbf{x}^{i},\mathbf{y}^{i})$ to compute toxicity score $\mathcal{T}^{i}$.
7:     while $\mathcal{T}^{i}>\mathcal{T}^{\text{max}}$ do
8:        Update $\mathcal{T}^{\text{max}}=\mathcal{T}^{i}$ and store current prompt $p^{i}$.
9:        Query $\mathcal{G}$ with $\mathbf{y}^{i}$ to generate a new prompt $p^{i}$.
10:        Generate new image $\mathbf{x}^{i}=D_{\theta}(p^{i})$.
11:        Query $\mathcal{M}$ with $(\mathbf{x}^{i},\mathbf{y}^{i})$ to compute updated toxicity score $\mathcal{T}^{i}$.
12:     end while
13:     Append the stored best prompt $p^{i}$ into $P$.
14:  end for
15:  for $step=1,...,N$ do
16:     Sample a batch of sentence-prompt pairs $\{(\mathbf{y}^{k},p^{k})\}_{k=1}^{b}$ from $Y$ and $P$.
17:     for each $(\mathbf{y}^{k},p^{k})$ in the batch do
18:        Generate image $\mathbf{x}^{k}=D_{\theta}(p^{k})$.
19:        Query $\mathcal{M}$ with $(\mathbf{x}^{k},\mathbf{y}^{k})$ to compute total reward $\mathcal{R}_{\text{total}}$.
20:     end for
21:     Update the Diffuser $D_{\theta}$ using the policy gradient: $\theta\leftarrow\theta+\alpha\cdot\nabla_{\theta}J$.
22:  end for
23:  return Fine-tuned Diffuser $D_{\theta}^{*}$.
5 Experiments
Table 1: Black-box attack results on the original attack set against LLaVA. Percentages (%) represent the rate of specific toxic attributes in the outputs, as evaluated by the Detoxify classifier and Perspective API. ‘Any*’ indicates the text exhibits at least one of the six toxic attributes.
Detoxify (%)
(%) Identity Attack Obscene Severe Toxicity Insult Threat Toxicity Any*
No Attack 5.34 39.69 0.76 19.84 0.00 58.01 58.01
Base Attack 3.05 38.16 0.00 21.37 1.52 55.72 55.72
Greedy attack 3.05 39.69 0.00 23.66 1.52 64.12 64.12
RTD w/o alignment 3.05 45.03 0.00 27.48 0.00 61.83 61.83
RTD w/ alignment (L) 3.81 44.27 1.52 24.42 1.52 62.59 62.59
RTD w/ alignment (M) 3.81 42.74 0.76 25.19 0.00 68.70 68.70
RTD w/ alignment (H) 3.05 40.45 0.00 25.19 0.76 63.35 63.35
Perspective API (%)
(%) Identity Attack Profanity Severe Toxicity Sexually Explicit Threat Toxicity Any*
No Attack 6.06 35.60 1.51 16.66 2.27 46.96 53.03
Base Attack 4.54 37.12 1.51 15.15 4.54 48.48 55.30
Greedy attack 4.54 40.90 1.51 16.66 3.78 52.27 58.33
RTD w/o alignment 4.54 42.42 3.78 21.96 0.75 54.54 58.33
RTD w/ alignment (L) 6.06 42.42 3.78 18.93 3.78 52.27 58.33
RTD w/ alignment (M) 5.30 41.66 1.51 20.45 3.03 56.81 61.36
RTD w/ alignment (H) 4.54 39.39 0.75 18.93 4.54 55.30 62.12
Datasets We use the RealToxicityPrompts benchmark Gehman et al. (2020), specifically its challenging subset of 1,199 text prompts designed to trigger toxic continuations Schick et al. (2021); Mehrabi et al. (2022); Qi et al. (2024). We select the first 132 prompts as our original attack set, as we observed that increasing the training prompt size barely improve generalization, but instead slowed down convergence. The remaining prompts are reserved as a hold-out set to evaluate cross-prompt transferability. While the original attack set is selected simply based on their ordering, subsequent analysis revealed that these prompts exhibit relatively lower baseline toxicity compared to the hold-out set. This characteristic makes them a more rigorous benchmark for evaluating RTD’s ability to induce harmful continuations. In particular, targeting prompts with initially less toxic outputs allows for a clearer demonstration of RTD’s effectiveness, as it highlights its capacity to generate harmful content even when baseline toxicity is minimal.
Metric: Toxicity Rate To evaluate the toxicity of continuations generated on the RealToxicityPrompts benchmark Gehman et al. (2020), we utilize the Detoxify classifier Hanu and Unitary team (2020) during training, and both Detoxify and the Perspective API during testing. Both classifiers assess toxicity across six predefined attributes. Each attribute’s toxicity score ranges from 0 (least toxic) to 1 (most toxic). The Perspective API is employed during testing to assess cross-classifier transferability. For each attribute, we calculate the proportion of continuations whose toxicity scores exceed a 0.5 threshold.
Implementation Details We use Stable Diffusion v1.5 Rombach et al. (2022) as the pre-trained diffusion model in our experiments. Gemini-1.5-Pro is employed to generate image prompts, while LLaVA-1.5-7B serves as the target VLM. Fine-tuning RTD is conducted using 8 NVIDIA A100 GPUs, where 6 are used for model training and 2 for reward computation. During reinforcement fine-tuning, we perform 600 policy update steps with a learning rate of 3e-4 and a batch size of 24 per GPU. We experiment with four different alignment reward weight $\lambda$ settings: 0 (w/o alignment), 0.05 (Low), 0.1 (Medium), and 0.2 (High).
5.1 Main Results
To evaluate the effectiveness of the Red Team Diffuser (RTD), we conduct black-box attacks on the original attack set and cross-prompt transfer attacks on the hold-out set against LLaVA. We compare several attack strategies: Base Attack: Gemini generates a single image prompt without optimization, providing a baseline for comparison. Greedy Attack: We iteratively refine the image prompt using the proposed greedy prompt search to maximize toxicity, selecting the most effective prompt at each step. RTD: We experiment with four alignment reward weight settings: 0 (w/o alignment), 0.05 (RTD w/ alignment (L)), 0.1 (RTD w/ alignment (M)), and 0.2 (RTD w/ alignment (H)), as discussed in the implementation details.
Table 2: Cross-prompt transfer attack results on the hold-out set against LLaVA. Percentages (%) represent the rate of specific toxic attributes in the outputs, as evaluated by the Detoxify classifier and Perspective API.
Detoxify (%)
(%) Identity Attack Obscene Severe Toxicity Insult Threat Toxicity Any*
No Attack 4.97 60.31 2.53 35.27 2.53 70.35 70.54
Base Attack 6.00 65.19 3.09 39.02 4.12 76.54 76.64
Greedy attack 5.44 66.22 3.00 40.15 3.47 77.39 77.57
RTD w/o alignment 6.00 67.44 3.09 42.02 4.22 79.36 79.45
RTD w/ alignment (L) 5.90 66.69 3.37 39.77 3.37 78.23 78.33
RTD w/ alignment (M) 5.53 66.32 3.28 40.52 4.69 78.42 78.70
RTD w/ alignment (H) 6.09 66.69 3.18 40.43 3.84 79.36 79.36
Perspective API (%)
(%) Identity Attack Profanity Severe Toxicity Sexually Explicit Threat Toxicity Any*
No Attack 6.74 61.38 5.43 21.36 3.84 68.22 70.00
Base Attack 7.87 65.04 5.99 22.68 5.90 73.19 75.63
Greedy attack 7.59 66.63 6.56 23.14 5.34 75.16 77.60
RTD w/o alignment 8.05 68.13 7.12 24.36 5.90 77.31 79.28
RTD w/ alignment (L) 8.15 67.29 7.12 23.89 5.62 75.35 77.50
RTD w/ alignment (M) 7.31 66.91 6.46 23.33 6.84 76.10 78.16
RTD w/ alignment (H) 7.68 66.91 6.65 23.43 5.99 76.47 78.81
Original Attack Set Table 1 presents the results on the original attack set against LLaVA, evaluated using both the Detoxify classifier and the Perspective API. The percentages represent the rate of specific toxic attributes in the outputs, with ‘Any*’ indicating whether the text exhibits at least one of the six toxic attributes. The proposed greedy prompt search demonstrates a significant increase in toxicity compared to the Base Attack. For example, using the Detoxify classifier, the Greedy Attack achieved an ‘Any*’ rate of 64.12%, compared to 55.72% for the Base Attack. This substantial improvement underscores the effectiveness of greedy prompt search in maximizing toxic content generation. Furthermore, the Red Team Diffuser (RTD) consistently outperforms the Greedy Attack, with RTD w/ alignment (M) achieving the highest ‘Any*’ rate of 68.70% according to Detoxify. This highlights the considerable impact of reinforced fine-tuning in enhancing the RTD’s ability to trigger toxic continuations. Although RTD was trained using Detoxify, the results from the Perspective API validate its cross-classifier transferability. For example, RTD w/ alignment (H) achieves the highest ‘Any*’ rate of 62.12% on Perspective API, demonstrating the robustness of RTD across different toxicity evaluation models.
Table 3: Cross-model transfer attack results on Gemini-1.5-flash. The percentages (%) represent the ‘Any*’ rates, as evaluated by the Detoxify classifier and Perspective API.
(%) Detoxify Perspective API
No Attack 76.23 78.26
RTD w/o alignment 80.30 81.49
RTD w/ alignment (L) 79.14 80.83
RTD w/ alignment (M) 81.33 82.52
RTD w/ alignment (H) 81.00 82.11
Hold-out Set Table 2 presents the results of cross-prompt transfer attacks on the hold-out set against LLaVA. Importantly, the prompts in the hold-out set were not seen during training, making this evaluation a strong indicator of the generalization capability of the proposed methods. Interestingly, the hold-out set exhibits a higher average baseline toxicity compared to the original attack set. This is likely due to the inherent nature of the hold-out prompts, which tend to be more inherently toxic and thus more prone to eliciting toxic responses from the target VLM. The Greedy Attack demonstrates a clear improvement over the Base Attack according to both the Detoxify classifier and the Perspective API. For instance, using Detoxify, the Greedy Attack achieved an ‘Any*’ rate of 77.57%, compared to 76.64% for the Base Attack. Similarly, with the Perspective API, the Greedy Attack achieved an ‘Any*’ rate of 77.60%, compared to 75.63% for the Base Attack. The Red Team Diffuser (RTD) consistently demonstrates superior performance across most metrics, achieving a peak ‘Any*’ rate of 79.45% according to Detoxify with RTD w/o alignment, as well as notable increases in specific attributes such as Obscene (67.44%) and Insult (42.02%). Furthermore, results from the Perspective API show that RTD w/o alignment achieves the highest ‘Any*’ rate of 79.28%, along with leading scores in attributes such as Profanity (68.13%) and Severe Toxicity (7.12%). These results clearly validate the effectiveness of our RTD in eliciting toxic continuations from diverse, unseen prompts. The consistent performance across both classifiers demonstrates its strong generalization capability beyond the training data.
Table 4: Cross-model transfer attack results on Llama-3.2-11B-Vision-Instruct. The percentages (%) represent the ‘Any*’ rates, as evaluated by the Detoxify classifier and Perspective API.
(%) Detoxify Perspective API
No Attack 15.02 18.68
RTD w/o alignment 41.01 43.90
RTD w/ alignment (L) 41.85 45.24
RTD w/ alignment (M) 41.81 44.37
RTD w/ alignment (H) 41.15 43.78
5.2 Cross-Model Transferability
We evaluate the cross-model transferability of the proposed Red Team Diffuser (RTD) by using the RTD trained on LLaVA Liu et al. (2024b). Transfer attacks are performed on two additional VLMs: Gemini-1.5-flash Team et al. (2023) and Llama-3.2-11B-Vision-Instruct. Tables 3 and 4 present the ‘Any*’ rates evaluated by both the Detoxify classifier and the Perspective API. These results demonstrate that RTD, trained specifically on a single VLM, can effectively trigger toxic content when applied to unseen models, underscoring its robustness and transferability across VLMs. Table 3 summarizes the attack performance on Gemini 1.5 Flash. RTD w/ alignment (M) achieves the highest ‘Any*’ rates of 81.33% according to Detoxify and 82.52% according to the Perspective API, significantly outperforming other configurations. This indicates that the RTD maintains strong transferability to Gemini-1.5-flash, even without further fine-tuning. Table 4 presents the results for Llama 3.2 11B Vision-Instruct, a VLM with more advanced alignment mechanisms. While the baseline (No Attack) exhibits relatively low toxicity rates (15.02% according to Detoxify and 18.68% according to Perspective API), the RTD significantly increases these metrics. Notably, the lower baseline toxicity of Llama 3.2 can be attributed to its high refusal rate for toxic continuation tasks, demonstrating a stronger alignment mechanism that often declines requests to generate potentially harmful content. In contrast, other VLMs, such as LLaVA and Gemini, rarely refuse these requests, leading to higher baseline toxicity levels. Despite this, RTD is able to effectively bypass Llama 3.2’s refusal mechanisms, significantly increasing toxicity rates. Specifically, RTD w/ alignment (L) achieves the highest ‘Any*’ rates of 41.85% according to Detoxify and 45.24% according to Perspective API, closely followed by RTD w/ alignment (M). These findings demonstrate that RTD effectively transfers its attack capabilities to VLMs with different architectures and scales, even when those models have stronger alignment mechanisms.
5.3 Visualizations
Figure 3 shows a qualitative comparison of images generated by the pre-trained Stable Diffusion (left) and the Red Team Diffuser (RTD) (right). For RTD-generated images, the alignment reward weights $\lambda$ progressively increase from left to right, starting with a weight of 0 in the leftmost column and followed by 0.05, 0.1, and 0.2. Despite being given the same image prompts, the outputs from the two models differ significantly. While the pre-trained Stable Diffusion produces neutral and benign visuals, RTD generates images that subtly evoke discomfort or unease. For instance, in the second row, RTD produces cartoonish and exaggerated caricatures that distort the features of an overweight person eating a sandwich, in contrast to the more neutral depiction by Stable Diffusion. In the third row, where the prompt involves generating an image of a woman, RTD creates visuals that are noticeably more sexually suggestive and explicit. Similarly, in the fourth row, RTD introduces additional elements, such as garbage and ominous shadows, into the school hallway scene generated by Stable Diffusion, thereby amplifying the potential for toxic continuations. The observed differences in visual elements and stylistic features illustrate how RTD effectively adapts image features to exploit vulnerabilities in VLMs, thereby enhancing its ability to induce harmful continuations.
Figure 3: Comparison of images generated by the pre-trained Stable Diffusion (left) and the Red Team Diffuser (right). Given the same image prompts, RTD produces images that subtly elicit discomfort or unease, making them more likely to trigger harmful continuations in VLMs.
6 Limitations
The RTD framework demonstrates strong performance in eliciting harmful outputs from VLMs, but several limitations deserve attention. First, this study focuses exclusively on toxic continuation tasks. Broader challenges, such as misinformation propagation and ethical biases in complex multimodal scenarios, remain unexplored. Second, RTD’s high computational demands, including the use of 8 NVIDIA A100 GPUs, may hinder its adoption by researchers with limited resources. Finally, RTD poses ethical risks, as its methods could potentially be misused by adversaries.
7 Conclusion
In this work, we propose the Red Team Diffuser (RTD) framework to identify and exploit vulnerabilities in VLMs, specifically in toxic continuation tasks. Leveraging greedy prompt search and reinforcement fine-tuning, RTD generates images that effectively trigger toxic continuations, exposing the limitations of existing alignment mechanisms. Experimental results validate the effectiveness of RTD, showing significant increases in toxicity rates across both black-box and cross-model settings. For instance, RTD raises LLaVA’s toxicity rates by 10.69% on the original attack set and 8.91% on a hold-out set, while also demonstrating strong cross-model transferability, with increases of 5.1% on Gemini and 26.83% on Llama-3.2-11B-Vision-Instruct. Qualitative analysis further highlights RTD’s ability to subtly manipulate visual features, generating images that are more provocative and capable of inducing harmful continuations. These findings highlight critical gaps in current alignment strategies and emphasize the need for more robust and comprehensive mechanisms to enhance the safety of VLMs in real-world applications.