### v2

# Red Team Diffuser: Exposing Toxic Continuation Vulnerabilities in Vision-Language Models via Reinforcement Learning

Abstract.
The growing deployment of large Vision-Language Models (VLMs) exposes critical safety gaps in their alignment mechanisms. While existing jailbreak studies primarily focus on VLMs’ susceptibility to harmful instructions, we reveal a fundamental yet overlooked vulnerability: toxic text continuation – where VLMs produce highly toxic completions when prompted with harmful text prefixes paired with semantically adversarial images. To systematically study this threat, we propose Red Team Diffuser (RTD), the first red teaming diffusion model that coordinates adversarial image generation and toxic continuation through reinforcement learning. Our key innovations include dynamic cross-modal attack and stealth-aware optimization. For toxic text prefixes from an LLM safety benchmark, we conduct greedy search to identify optimal image prompts that maximally induce toxic completions. The discovered image prompts then drive RL-based diffusion model fine-tuning, producing semantically-aligned adversarial images that boost toxicity rates. Stealth-aware optimization introduces joint adversarial rewards that balance toxicity maximization (via Detoxify classifier) and stealthiness (via BERTScore), circumventing traditional noise-based adversarial patterns. Experimental results demonstrate the effectiveness of RTD, increasing the toxicity rate of LLaVA outputs by 10.69% over text-only baselines on the original attack set and 8.91% on unseen set, proving generalization capability. Moreover, RTD exhibits strong cross-model transferability, raising the toxicity rate by 5.1% on Gemini and 26.83% on LLaMA. Our findings expose two critical flaws in current VLM alignment: (1) failure to prevent toxic continuation from harmful prefixes, and (2) overlooking cross-modal attack vectors. These results necessitate a paradigm shift toward multimodal red teaming in safety evaluations. We will publicly release the RTD framework to facilitate future research on robust VLM alignment.
Disclaimer: This paper contains content that may be disturbing or offensive.
Red Teaming, Vision-Language Models, Diffusion Models, Reinforcement Learning
ccs: Computing methodologies Artificial intelligence
1. Introduction
Figure 1. Example of a successful jailbreak on LLaVA. The proposed Red Team Diffuser (RTD) generates toxic-triggering images despite using the same prompt as general-purpose diffusion model (Rombach et al., 2022).
The rapid deployment of large Vision-Language Models (VLMs), such as GPT-4o (Achiam et al., 2023) and Gemini (Team et al., 2023), has exposed critical blind spots in their multimodal alignment mechanisms. While existing safety strategies predominantly focus on preventing responses to explicit harmful instructions (e.g., ”How to build a bomb?”) through techniques like Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022), they often fail to address the risks posed by toxic text continuation tasks - where VLMs amplify harmful content when completing partially toxic text prompts paired with semantically adversarial images (Figure 1). This vulnerability originates from two research gaps: (a) Current red teaming methods focus on explicit instruction attacks rather than contextual toxicity induction. (b) Existing safety protocols neglect cross-modal attack vectors where images amplify output toxicity.
The multimodal nature of VLMs introduces unique safety challenges: identical text inputs paired with different images can yield drastically different toxicity levels (Liu et al., 2023), sometimes in unintuitive ways. For instance, Li et al. (Li et al., 2024) demonstrated that pairing jailbreak prompts with even a blank image can significantly increase attack success rates. However, systematically identifying images that effectively trigger toxic continuations remains an open and challenging problem. Recent progress in white-box jailbreak techniques (Qi et al., 2024; Wang et al., 2024b) has explored the use of adversarial images optimized directly against model parameters. While powerful, such gradient-based approaches are impractical in real-world scenarios, particularly within black-box VLM-as-a-Service contexts where full model access is unavailable. Additionally, adversarial images generated through gradient-based methods typically exhibit noticeable distortions, making them susceptible to detection and defense mechanisms (Nie et al., 2022; Zhang et al., 2023; Zhao et al., 2024). Moreover, studies have demonstrated that these adversarial images typically show limited transferability between VLMs, even after successfully jailbreaking multiple VLMs simultaneously (Schaeffer et al., [n. d.]).
These limitations motivate us to develop stealthy black-box attacks that produce natural adversarial images without full model access. Existing black-box jailbreak methods typically adopt suboptimal paradigms, rely on manually engineered pipelines involving typographic manipulations or pre-trained diffusion models to generate query-relevant images (Gong et al., 2023; Liu et al., 2023; Teng et al., 2024). However, such methods have critical drawbacks. First, structure-based attacks inherently constrain the attack surface, resulting in predictable patterns that are easily mitigated by modern defense mechanisms (Wang et al., 2024a) and detectable by OCR systems. Second, such images are primarily designed to bypass sensitive text detection, rather than actively inducing harmful continuations, limiting their applicability to implicit toxicity tasks.
More recently, automated red-teaming frameworks have emerged to systematically uncover vulnerabilities in VLMs. For instance, IDEATOR (Wang et al., 2024c) introduced the ”VLMs against VLMs” paradigm, employing a red-team VLM to generate jailbreak prompts paired with images from general-purpose diffusion models (Rombach et al., 2022). While innovative, this approach inherits a fundamental limitation: the inherent conflict between diffusion models’ original training objective (faithful visual-textual alignment) and red teaming goals (toxicity maximization). Off-the-shelf diffusers like Stable Diffusion (Rombach et al., 2022) are optimized to avoid generating harmful content during pretraining, creating an intrinsic bias against jailbreaking effectiveness.
This highlights a critical yet unexplored research opportunity: Can we reconfigure diffusion models to actively induce toxicity while maintaining semantic plausibility? Current approaches treat image generation and toxicity induction as decoupled processes - first generating ”safe” images with pretrained diffusers, then hoping they coincidentally trigger harmful outputs. Our work overturns this passive paradigm through reinforced diffusion alignment, where the diffusion model is actively tuned to: (1) maximize toxicity through RL-driven reward signals; (2) preserve semantic coherence with image prompts; (3) dynamically adapt to target VLM’s defense patterns. By transforming diffusion models from general-purpose generators into active toxicity amplifiers, we establish a new paradigm for multimodal red teaming that fundamentally rethinks how visual modalities can be weaponized against VLMs.
In this work, we propose Red Team Diffuser (RTD), a novel framework that redefines multimodal jailbreak through semantically adversarial image synthesis. Departing from conventional static attack patterns, RTD introduces a dynamic two-phase optimization paradigm. First, we leverage the contextual reasoning capabilities of an LLM to perform a greedy search over image prompts, identifying those that maximize the toxicity of VLM responses and are thus highly effective for red teaming. Second, the discovered prompts drive RL-based fine-tuning of a diffusion model, using a dual-objective reward: a toxicity reward to amplify the harmfulness of the induced outputs, and an image-prompt alignment reward to ensure semantic consistency as a regularization term. This transforms standard diffusion models into toxicity-aware generators, fundamentally altering their output distribution toward vulnerable regions of target VLMs’ multimodal alignment space.
The main contributions of our work are threefold:
  • New Attack Paradigm: We propose Red Team Diffuser (RTD), the first framework to integrate reinforcement learning with diffusion models for semantically adversarial image generation, establishing a new paradigm for multimodal red teaming. Unlike prior approaches that rely on static patterns or unmodified diffusers, RTD actively co-optimizes both image prompts and diffusers for toxicity induction.
  • Technical Innovations: We employ an LLM to automatically generate diverse image prompts, followed by iterative toxicity evaluation to identify the most effective triggers. Furthermore, we introduce an RL-based fine-tuning strategy with dual rewards, balancing toxicity maximization and semantic preservation, thereby ensuring both high attack potency and natural image quality.
  • Empirical Contributions: We demonstrate a 10.69% increase in toxicity rate over text-only attacks on LLaVA. Furthermore, RTD exhibits strong transferability across prompts and models, achieving an 8.91% increase on a hold-out set, along with toxicity rate gains of 5.1% on Google’s Gemini and 26.83% on Meta’s LLaMA. Our findings reveal a critical vulnerability in cross-modal toxicity amplification, where semantically adversarial images exacerbate text-based harm.
2. Related Work
2.1. Large Vision-Language Models
Large Vision-Language Models (VLMs) extend Large Language Models (LLMs) by integrating visual modalities through an image encoder and a cross-modal alignment module. This design enables VLMs to perform instruction-following and open-ended reasoning over both text and images. Recent open-source VLMs have explored various architecture designs. LLaVA (Liu et al., 2024a) connects a CLIP encoder (Radford et al., 2021) with LLaMA (Touvron et al., 2023), using GPT-4 (Achiam et al., 2023) to synthesize multimodal instruction-following data for end-to-end fine-tuning. MiniGPT-4 (Zhu et al., 2023) employs a linear projection to align a frozen ViT (Dosovitskiy et al., 2020) with Vicuna (Chiang et al., 2023), using pretraining on large-scale image-text pairs followed by high-quality fine-tuning. InstructBLIP (Dai et al., 2023), based on BLIP-2 (Li et al., 2023b), introduces an instruction-aware Query Transformer and reformats 26 public datasets to enhance instruction understanding. While these models achieve impressive performance, the incorporation of visual inputs significantly broadens the attack surface compared to text-only LLMs. Recent studies show that visual perturbations and cross-modal inconsistencies can be exploited to bypass safety mechanisms or elicit harmful behavior (Shayegani et al., 2023b; Liu et al., 2024b; Li et al., 2023a; Ma et al., 2025). These findings highlight the urgent need to evaluate VLM robustness under multimodal threats—an area this work directly addresses.
Figure 2. Overview of the proposed method. Given an incomplete sentence, Gemini is used to generate an image prompt designed to maximize the likelihood of toxic continuation. The diffusion model generates an image based on this prompt. Toxicity and alignment rewards are then computed from the target VLM (LLaVA) response and used to reinforce fine-tuning of the diffusion model.
2.2. Jailbreak Attacks against VLMs
Jailbreak attacks against VLMs exploit the interaction between visual and textual inputs to bypass safeguards, enabling the generation of harmful content. These attacks can be categorized as either white-box or black-box methods. White-box jailbreak attacks (Bagdasaryan et al., 2023; Bailey et al., 2023; Shayegani et al., 2023a; Carlini et al., 2024; Qi et al., 2024; Niu et al., 2024; Wang et al., 2024b) typically leverage gradient information to perturb input images or text to increase the likelihood of harmful outputs. While effective, they require full or partial access to model parameters, making them less practical and more easily detectable (Nie et al., 2022; Zhang et al., 2023). In contrast, black-box jailbreak attacks present more practical and scalable solutions, as they do not require direct access to model parameters but instead exploit external vulnerabilities. Despite their advantages, many existing black-box approaches rely heavily on typography-based manipulations, making them susceptible to detection techniques like OCR. For example, FigStep (Gong et al., 2023) converts harmful text into images using typography to bypass safety mechanisms. Similarly, approaches like MM-SafetyBench and VRP (Liu et al., 2023; Ma et al., 2024) also employ typography-driven attacks as part of their methods, compromising their stealthiness against advanced detection systems. To address these limitations, the concept of red team VLMs has progressed with IDEATOR (Wang et al., 2024c), which autonomously generates jailbreak image-text pairs in a black-box setting. By offering a scalable and automated pipeline for discovering jailbreak prompts, IDEATOR demonstrates the promise of using VLMs themselves to expose alignment weaknesses. While IDEATOR focuses on prompt-level red teaming, our work introduces a fundamentally different and complementary attack paradigm: reinforced diffusion-based image optimization. Rather than relying on VLM-generated inputs, we investigate how semantically adversarial images—crafted via reinforcement learning—can effectively induce toxic text continuations when paired with fixed text prefixes. This setup presents a significantly more challenging threat model: it removes reliance on typographic or prompt engineering tricks and explores the vulnerability of VLMs to unrestricted, natural-looking visual stimuli. By doing so, we establish a rigorous framework for evaluating VLM robustness against cross-modal attacks, highlighting a critical blind spot in existing alignment strategies.
3. Preliminary
Denoising Diffusion Policy Optimization  RTD employs the Denoising Diffusion Policy Optimization (DDPO) algorithm (Black et al., 2023) to fine-tune a pretrained diffusion model (Rombach et al., 2022). Unlike traditional supervised learning with labeled data, DDPO leverages reinforcement learning to optimize the model by framing the iterative denoising process as a multi-step Markov Decision Process (MDP). At each step, actions (denoising predictions) are chosen to maximize the cumulative reward. The state of the MDP at time step $t$ is defined as $\mathbf{s}_{t}=\left(\mathbf{c},t,\mathbf{x}_{T-t}\right)$, where $\mathbf{c}$ represents the context, $t$ is the current time step, $\mathbf{x}_{T-t}$ is the noisy sample at time step $t$, and $T$ is the total number of denoising steps. Note that the MDP is a denoising process, i.e., the reverse diffusion process that recreates the true sample $\mathbf{x}_{0}$ from a Gaussian noise input $\mathbf{x}_{T}\sim\mathcal{N}(0,1)$. The agent (i.e., the diffuser)’s action at time step $t$ is defined as $a_{t}=\mathbf{x}_{T-t-1}$, which corresponds to the predicted previous state (i.e., the denoised sample) in the reverse process. The agent’s policy is thus $\pi_{\theta}\left(\mathbf{a}_{t}\mid\mathbf{s}_{t}\right)=p_{\theta}\left(\mathbf{x}_{T-t-1}\mid\mathbf{x}_{T-t},\mathbf{c}\right)$. The task reward for the agent $R(\mathbf{s}_{t},\mathbf{a}_{t})$ is structured as $R(\mathbf{s}_{t},\mathbf{a}_{t})=r(\mathbf{x}_{0},\mathbf{c})\text{ if }t=T\text{ else }0$. That is, the reward is non-zero only at the final time step ($t=T$), and the cumulative reward of a trajectory is equal to $r(\mathbf{x}_{0},\mathbf{c})$.
The goal of DDPO is to maximize the expected cumulative reward, which is defined over the sample distribution $p_{\theta}(x_{0}\mid c)$ and the context distribution $p(c)$:
$\mathcal{J}(\theta)=\mathbb{E}_{\mathbf{c}\sim p(\mathbf{c}),\mathbf{x}_{T}\sim\mathcal{N}(0,1),a\sim\pi_{\theta}(\cdot|s)}\left[R\left(s,a\right)\right].$
This objective encourages the model to learn a policy that generates samples close to the target distribution. To optimize the objective across multiple steps, we employ an importance sampling-based policy gradient estimator (Kakade and Langford, 2002):
$\nabla_{\theta}\mathcal{J}=\mathbb{E}\Bigg{[}\sum_{t=0}^{T}\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})}\nabla_{\theta}\log\pi_{\theta}(a_{t}|s_{t})R\left(s,a\right)\Bigg{]},$
where the expectation is taken over denoising trajectories generated by the previous policy $\pi_{\theta_{\text{old}}}$. This estimator leverages the ratio of the current model’s likelihood to the previous model’s likelihood, allowing for more efficient updates.
4. Methodology
4.1. Threat Model
Attack Goals  We focus on the sentence continuation task, an underexplored area in alignment mechanisms that poses unique challenges due to the open-ended nature of model outputs. The attacker aims to bypass the model’s safety measures and induce toxic continuations, such as the generation of obscene or profane language. Toxic continuations can also be viewed as a form of multimodal prompt injection, where carefully designed text prefixes and image triggers manipulate the model’s output into generating harmful content.
Adversary Capabilities  We assume the attacker has only black-box access to the victim VLM, simulating real-world interactions via external interfaces. By reinforcement fine-tuning red team diffusers based on the VLM’s feedback, RTD learns the VLM’s behavior patterns and identifies vulnerabilities, refining its policy to trigger toxic content.
4.2. Greedy Prompt Search
The effectiveness of our toxic continuation attacks hinges on identifying image prompts that maximally amplify toxicity in VLM-generated outputs. To overcome the limitations of directly using LLM-generated prompts, we introduce a dynamic, feedback-driven search procedure designed to discover high-risk image prompts tailored to specific toxic text prefixes. We employ Google’s Gemini Pro (Team et al., 2023)—with safety settings disabled—as a red team LLM to propose candidate image prompts conditioned on each text prefix from the LLM safety benchmark. Each image prompt is required to retain the semantic context of the prefix while subtly amplifying its harmful intent, thereby increasing the likelihood of inducing toxic completions. To enhance control over prompt generation, we apply in-context learning (Brown, 2020) by providing exemplars of previously identified harmful sentence continuations in the system prompt, which serve as implicit guidance for Gemini to produce image prompts consistent with known harmful patterns.
Despite Gemini’s capabilities, its generated prompts may still be suboptimal in terms of triggering maximal toxicity. To this end, we propose a greedy prompt search algorithm to iteratively refine image prompts based on observed toxicity feedback. At each iteration, Gemini generates a new candidate prompt, which is used to synthesize an image via the diffusion model. This image, paired with the fixed text prefix, is input to the target VLM, and the resulting continuation is scored using a toxicity classifier (e.g., Detoxify (Hanu and Unitary team, 2020)). If the toxicity score increases relative to the previous iteration, the new image prompt is retained and the search continues; otherwise, the process terminates. This loop yields near-optimal image prompts that serve as strong initialization points for downstream adversarial image generation.
4.3. Reward Functions for RTD
A contribution of RTD lies in its dual-reward optimization, which explicitly balances attack effectiveness and semantic plausibility. We introduce two complementary rewards: a toxicity reward that drives harmful continuation, and an alignment reward that ensures prompt-consistent image generation.
Toxicity Reward: In our toxic continuation task, we prompt the target VLM $\mathcal{M}$ to extend a given sentence $\mathbf{y}$, paired with an image $\mathbf{x}$ generated by the RTD. The toxicity of the generated continuation is then assessed using the Detoxify classifier (Hanu and Unitary team, 2020), which evaluates six toxicity-related attributes (e.g., insult, threat, identity attack), each scored between 0 (least toxic) and 1 (most toxic). We use the average of these scores as the toxicity reward to guide diffusion-based adversarial image synthesis:
$\mathcal{R}_{\text{tox}}(\mathcal{M}(\mathbf{x},\mathbf{y}))=\frac{1}{6}\sum_{i=1}^{6}\text{toxicity}_{i}(\mathcal{M}(\mathbf{x},\text{``Please}\\ \text{``extend the following sentence: "}+\mathbf{y}))$
Alignment Reward: To encourage semantic alignment between the input prompt $p$ and the VLM’s interpretation of the generated image, we introduce an alignment reward that regularizes the image generation process. Following (Black et al., 2023), we use BERTScore (Zhang et al., 2019) to measure the similarity between the image prompt and the caption produced by the victim VLM. Specifically, the alignment reward is defined as:
$\mathcal{R}_{\text{align}}(p,\mathcal{M}(\mathbf{x}))=\frac{1}{|p|}\sum_{i=1}^{|p|}\max_{j}\big{(}S(p_{i},\mathcal{M}(\mathbf{x},\\ \text{``Describe the image in short."})_{j})\big{)}$
[…]
Total Reward: The total reward is formulated as a weighted sum of the toxicity reward and the alignment reward, controlled by a hyperparameter $\lambda$:
$\mathcal{R}_{\text{total}}=\mathcal{R}_{\text{tox}}+\lambda\mathcal{R}_{\text{align}}$
This combined reward allows the RTD to generate both toxic-inducing and semantically aligned image, thus achieving a balance between maximizing the harmfulness and maintaining image coherence. The overview of RTD is illustrated in Figure 2, while the detailed strategy is summarized in Algorithm 1.
Algorithm 1 Training Procedure of Red Team Diffuser
Require: Pre-trained Diffuser $D_{\theta}$, frozen LLM (e.g., Gemini) $\mathcal{G}$, target VLM $\mathcal{M}$, sentence continuation corpus $Y:=\{\mathbf{y}^{i}\}_{i=1}^{m}$, batch size $b$, learning rate $\alpha$, policy update steps $N$;
1:  Initialize $D_{\theta}$ with pre-trained weights and $P=\{\}$.
2:  for each sentence $\mathbf{y}^{i}\in Y$ do
3:     Initialize toxicity score $\mathcal{T}^{\text{max}}=0$.
4:     Query $\mathcal{G}$ with $\mathbf{y}^{i}$ to generate initial image prompt $p^{i}$.
5:     Generate initial image $\mathbf{x}^{i}=D_{\theta}(p^{i})$.
6:     Query $\mathcal{M}$ with $(\mathbf{x}^{i},\mathbf{y}^{i})$ to compute toxicity score $\mathcal{T}^{i}$.
7:     while $\mathcal{T}^{i}>\mathcal{T}^{\text{max}}$ do
8:        Update $\mathcal{T}^{\text{max}}=\mathcal{T}^{i}$ and store current prompt $p^{i}$.
9:        Query $\mathcal{G}$ with $\mathbf{y}^{i}$ to generate a new prompt $p^{i}$.
10:        Generate new image $\mathbf{x}^{i}=D_{\theta}(p^{i})$.
11:        Query $\mathcal{M}$ with $(\mathbf{x}^{i},\mathbf{y}^{i})$ to compute updated toxicity score $\mathcal{T}^{i}$.
12:     end while
13:     Append the stored best prompt $p^{i}$ into $P$.
14:  end for
15:  for $step=1,...,N$ do
16:     Sample a batch of sentence-prompt pairs $\{(\mathbf{y}^{k},p^{k})\}_{k=1}^{b}$ from $Y$ and $P$.
17:     for each $(\mathbf{y}^{k},p^{k})$ in the batch do
18:        Generate image $\mathbf{x}^{k}=D_{\theta}(p^{k})$.
19:        Query $\mathcal{M}$ with $(\mathbf{x}^{k},\mathbf{y}^{k})$ to compute total reward $\mathcal{R}_{\text{total}}$.
20:     end for
21:     Update the Diffuser $D_{\theta}$ using the policy gradient: $\theta\leftarrow\theta+\alpha\cdot\nabla_{\theta}J$.
22:  end for
23:  return Fine-tuned Diffuser $D_{\theta}^{*}$.
5. Experiments
Datasets We use the RealToxicityPrompts benchmark (Gehman et al., 2020), specifically its challenging subset of 1,199 text prompts designed to trigger toxic continuations (Schick et al., 2021; Mehrabi et al., 2022; Qi et al., 2024). We select the first 132 prompts as our original attack set, as we observed that increasing the training prompt size barely improve generalization, but instead slowed down convergence. The remaining prompts are reserved as a hold-out set to evaluate cross-prompt transferability. While the original attack set is selected simply based on their ordering, subsequent analysis revealed that these prompts exhibit relatively lower baseline toxicity compared to the hold-out set. This characteristic makes them a more rigorous benchmark for evaluating RTD’s ability to induce harmful continuations. In particular, targeting prompts with initially less toxic outputs allows for a clearer demonstration of RTD’s effectiveness, as it highlights its capacity to generate harmful content even when baseline toxicity is minimal.
Metric: (1) Toxicity Rate. To evaluate the toxicity of continuations generated on the RealToxicityPrompts benchmark (Gehman et al., 2020), we use the Detoxify classifier (Hanu and Unitary team, 2020) during training, and both Detoxify and the Perspective API during testing. These classifiers assess toxicity across six predefined attributes, with scores ranging from 0 (least toxic) to 1 (most toxic). For evaluation, we compute the proportion of continuations whose toxicity score exceeds a threshold of 0.5 for each attribute. The use of the Perspective API at test time allows us to assess cross-classifier generalization and mitigate risks of reward hacking or classifier-specific bias. (2) CLIPScore. To evaluate the semantic alignment between generated images and their prompts, we use CLIPScore (Hessel et al., 2021), a reference-free metric based on the cosine similarity between CLIP-encoded image and text features. Following the original formulation: $\text{CLIPScore}=\alpha\cdot\max(0,\cos(f_{\text{image}},f_{\text{text}}))$, where $f_{\text{image}}$ and $f_{\text{text}}$ denote CLIP embeddings, and $\alpha$ is a scaling factor (set to 2.5 following (Hessel et al., 2021)) for alignment with human judgment. In our setting, CLIPScore serves as an auxiliary metric to verify that adversarial images preserve semantic plausibility without compromising visual coherence.
Implementation Details We use Stable Diffusion v1.5 (Rombach et al., 2022) as the pre-trained diffusion model in our experiments. Gemini-1.5-Pro is employed to generate image prompts, while LLaVA-1.5-7B serves as the target VLM. Fine-tuning RTD is conducted using 8 NVIDIA A100 GPUs, where 6 are used for model training and 2 for reward computation. During reinforcement fine-tuning, we perform 600 policy update steps with a learning rate of 3e-4 and a batch size of 24 per GPU. We experiment with four different alignment reward weight $\lambda$ settings: 0 (w/o alignment), 0.05 (Low), 0.1 (Medium), and 0.2 (High).
5.1. Main Results
To evaluate the effectiveness of Red Team Diffuser (RTD), we conduct black-box toxic continuation attacks on the original attack set and cross-prompt transfer attacks on a hold-out set using LLaVA as the target VLM. We compare RTD against several baseline strategies to assess its advantage in inducing harmful outputs. We first include a Text-only Attack, in which the model is prompted solely with toxic text prefixes without any accompanying image. This serves as a unimodal baseline to quantify the additional effect of visual inputs on toxicity. Next, we define the Single-Prompt Attack, where Gemini generates a single image prompt per prefix without any optimization. This provides a naive multimodal baseline, capturing the effect of directly pairing a non-optimized image with a toxic prefix. The Greedy-Prompt Attack improves upon this by iteratively refining the image prompt using our proposed greedy search algorithm. Finally, we evaluate the full RTD framework, which incorporates reinforcement learning to fine-tune the diffusion model guided by toxicity and alignment rewards. We experiment with four alignment reward weight settings: 0 (w/o alignment), 0.05 (RTD w/ alignment (L)), 0.1 (RTD w/ alignment (M)), and 0.2 (RTD w/ alignment (H)), as discussed in the implementation details.
Table 1. Black-box attack results on the original attack set against LLaVA. Percentages (%) represent the rate of specific toxic attributes in the outputs, as evaluated by the Detoxify classifier and Perspective API. ‘Any*’ indicates the text exhibits at least one of the six toxic attributes.
Detoxify (%)
(%) Identity Attack Obscene Severe Toxicity Insult Threat Toxicity Any*
Text-only Attack 5.34 39.69 0.76 19.84 0.00 58.01 58.01
Single-Prompt Attack 3.05 38.16 0.00 21.37 1.52 55.72 55.72
Greedy-Prompt Attack 3.05 39.69 0.00 23.66 1.52 64.12 64.12
RTD w/o alignment 3.05 45.03 0.00 27.48 0.00 61.83 61.83
RTD w/ alignment (L) 3.81 44.27 1.52 24.42 1.52 62.59 62.59
RTD w/ alignment (M) 3.81 42.74 0.76 25.19 0.00 68.70 68.70
RTD w/ alignment (H) 3.05 40.45 0.00 25.19 0.76 63.35 63.35
Perspective API (%)
(%) Identity Attack Profanity Severe Toxicity Sexually Explicit Threat Toxicity Any*
Text-only Attack 6.06 35.60 1.51 16.66 2.27 46.96 53.03
Single-Prompt Attack 4.54 37.12 1.51 15.15 4.54 48.48 55.30
Greedy-Prompt Attack 4.54 40.90 1.51 16.66 3.78 52.27 58.33
RTD w/o alignment 4.54 42.42 3.78 21.96 0.75 54.54 58.33
RTD w/ alignment (L) 6.06 42.42 3.78 18.93 3.78 52.27 58.33
RTD w/ alignment (M) 5.30 41.66 1.51 20.45 3.03 56.81 61.36
RTD w/ alignment (H) 4.54 39.39 0.75 18.93 4.54 55.30 62.12
Table 2. Cross-prompt transfer attack results on the hold-out set against LLaVA. Percentages (%) represent the rate of specific toxic attributes in the outputs, as evaluated by the Detoxify classifier and Perspective API.
Detoxify (%)
(%) Identity Attack Obscene Severe Toxicity Insult Threat Toxicity Any*
Text-only Attack 4.97 60.31 2.53 35.27 2.53 70.35 70.54
Single-Prompt Attack 6.00 65.19 3.09 39.02 4.12 76.54 76.64
Greedy-Prompt Attack 5.44 66.22 3.00 40.15 3.47 77.39 77.57
RTD w/o alignment 6.00 67.44 3.09 42.02 4.22 79.36 79.45
RTD w/ alignment (L) 5.90 66.69 3.37 39.77 3.37 78.23 78.33
RTD w/ alignment (M) 5.53 66.32 3.28 40.52 4.69 78.42 78.70
RTD w/ alignment (H) 6.09 66.69 3.18 40.43 3.84 79.36 79.36
Perspective API (%)
(%) Identity Attack Profanity Severe Toxicity Sexually Explicit Threat Toxicity Any*
Text-only Attack 6.74 61.38 5.43 21.36 3.84 68.22 70.00
Single-Prompt Attack 7.87 65.04 5.99 22.68 5.90 73.19 75.63
Greedy-Prompt Attack 7.59 66.63 6.56 23.14 5.34 75.16 77.60
RTD w/o alignment 8.05 68.13 7.12 24.36 5.90 77.31 79.28
RTD w/ alignment (L) 8.15 67.29 7.12 23.89 5.62 75.35 77.50
RTD w/ alignment (M) 7.31 66.91 6.46 23.33 6.84 76.10 78.16
RTD w/ alignment (H) 7.68 66.91 6.65 23.43 5.99 76.47 78.81
Original Attack Set Table 1 presents the black-box attack results on the original attack set against LLaVA, evaluated using both the Detoxify classifier and the Perspective API. The percentages represent the rate of specific toxic attributes in the generated outputs, with ‘Any*’ indicating whether a continuation exhibits at least one of the six toxic attributes. We begin by examining the Text-only Attack, where only the toxic text prefix is provided without any visual input. Surprisingly, LLaVA still produces toxic continuations at a non-trivial rate—achieving an ‘Any*’ rate of 58.01% under Detoxify and 53.03% under Perspective. This indicates that current alignment mechanisms in VLMs remain insufficient for preventing harmful completions even in unimodal settings, exposing a fundamental vulnerability in their default safety alignment. When introducing visual input, the Single-Prompt Attack slightly alters the toxicity pattern but does not significantly improve attack effectiveness, serving as a weak multimodal baseline. In contrast, the Greedy-Prompt Attack significantly amplifies toxicity. For instance, using the Detoxify classifier, the ‘Any*’ rate increases from 55.72% (Single-Prompt) to 64.12%, illustrating the benefit of greedy prompt search in finding high-risk image prompts. The full Red Team Diffuser (RTD) further enhances attack strength. RTD w/ alignment (M) achieves the highest ‘Any*’ rate of 68.70% on Detoxify, indicating that reinforced fine-tuning of the diffusion model allows for stronger and more consistent toxic triggers. Notably, RTD also demonstrates strong cross-classifier transferability: RTD w/ alignment (H) achieves the highest ‘Any*’ rate of 62.12% under the Perspective API, despite being trained only with Detoxify. This confirms that the toxic behaviors induced by RTD are not artifacts of a single classifier but reflect a general vulnerability in VLM alignment under multimodal inputs.
Table 3. Cross-model transfer attack results on Gemini-1.5-flash. Percentages (%) represent the rate of specific toxic attributes in the outputs, as evaluated by the Detoxify classifier and Perspective API.
Detoxify (%)
(%) Identity Attack Obscene Severe Toxicity Insult Threat Toxicity Any*
Text-only Attack 4.18 61.75 2.25 34.64 2.84 76.23 76.23
RTD w/o alignment 4.94 65.29 2.68 38.89 4.19 80.13 80.30
RTD w/ alignment (L) 5.36 64.40 3.09 39.69 4.69 79.06 79.14
RTD w/ alignment (M) 5.27 65.77 3.17 39.74 4.43 81.25 81.33
RTD w/ alignment (H) 5.54 66.63 3.52 39.91 4.45 80.84 81.00
Perspective API (%)
(%) Identity Attack Profanity Severe Toxicity Sexually Explicit Threat Toxicity Any*
Text-only Attack 5.76 60.95 3.17 19.14 5.01 74.41 78.26
RTD w/o alignment 7.20 63.56 4.43 22.02 6.86 77.97 81.49
RTD w/ alignment (L) 7.19 63.01 4.68 21.50 6.35 76.98 80.83
RTD w/ alignment (M) 7.02 64.29 4.68 22.57 7.52 79.01 82.52
RTD w/ alignment (H) 7.38 64.73 4.86 22.16 7.13 79.09 82.11
Hold-out Set Table 2 presents the results of cross-prompt transfer attacks on the hold-out set against LLaVA. Since these prompts were not seen during training, this evaluation provides a robust measure of the generalization ability of our proposed methods. Interestingly, the hold-out set exhibits a higher average baseline toxicity compared to the original attack set. This is likely due to the inherent nature of the hold-out prompts, which tend to be more inherently toxic and thus more prone to eliciting toxic responses from the target VLM. The Greedy-Prompt Attack demonstrates a clear improvement over the Single-Prompt Attack according to both the Detoxify classifier and the Perspective API. For instance, using Detoxify, the Greedy-Prompt Attack achieved an ‘Any*’ rate of 77.57%, compared to 76.64% for the Single-Prompt Attack. Similarly, with the Perspective API, the Greedy-Prompt Attack achieved an ‘Any*’ rate of 77.60%, compared to 75.63% for the Single-Prompt Attack. More notably, the full Red Team Diffuser (RTD) consistently demonstrates superior performance across most metrics, achieving a peak ‘Any*’ rate of 79.45% according to Detoxify with RTD w/o alignment, as well as notable increases in specific attributes such as Obscene (67.44%) and Insult (42.02%). Furthermore, results from the Perspective API show that RTD w/o alignment achieves the highest ‘Any*’ rate of 79.28%, along with leading scores in attributes such as Profanity (68.13%) and Severe Toxicity (7.12%). These consistent gains clearly validate the strong generalization capability of RTD across diverse, unseen prompts and across different toxicity classifiers, underscoring its effectiveness as a black-box multimodal jailbreak method.
Table 4. Cross-model transfer attack results on Llama-3.2-11B-Vision-Instruct. Percentages (%) represent the rate of specific toxic attributes in the outputs, as evaluated by the Detoxify classifier and Perspective API.
Detoxify (%)
(%) Identity Attack Obscene Severe Toxicity Insult Threat Toxicity Any*
Text-only Attack 0.25 11.18 0.00 3.92 0.16 14.52 15.02
RTD w/o alignment 1.58 32.49 0.16 16.54 2.84 39.93 41.01
RTD w/ alignment (L) 1.67 32.83 0.33 15.37 2.50 41.43 41.85
RTD w/ alignment (M) 1.58 32.63 0.33 16.77 2.25 41.23 41.81
RTD w/ alignment (H) 2.17 31.30 0.16 16.27 2.25 40.56 41.15
Perspective API (%)
(%) Identity Attack Profanity Severe Toxicity Sexually Explicit Threat Toxicity Any*
Text-only Attack 0.50 13.84 0.00 3.16 0.66 15.59 18.68
RTD w/o alignment 2.67 35.47 1.16 7.26 2.83 39.23 43.90
RTD w/ alignment (L) 2.58 35.30 1.00 8.01 2.58 40.98 45.24
RTD w/ alignment (M) 3.25 34.27 1.41 7.92 3.16 39.69 44.37
RTD w/ alignment (H) 3.58 34.19 0.58 8.34 2.83 39.94 43.78
5.2. Cross-Model Transferability
We evaluate the cross-model transferability of our proposed method by using the RTD trained solely on LLaVA (Liu et al., 2024a). Transfer attacks are performed on two unseen VLMs: Gemini-1.5-flash (Team et al., 2023) and Llama-3.2-11B-Vision-Instruct. Tables 3 and 4 present the toxicity rates evaluated by both the Detoxify classifier and the Perspective API. These results demonstrate that RTD generalizes effectively across models, highlighting its robustness and potential threat as a transferable multimodal jailbreak method.
Table 3 presents the results on Gemini-1.5-flash. Despite being a commercial VLM, Gemini exhibits high baseline toxicity under the Text-only Attack, with ‘Any*’ rates reaching 76.23% (Detoxify) and 78.26% (Perspective), indicating that harmful completions are already prevalent even without adversarial visual input. Furthermore, all RTD variants elevate toxicity levels, with RTD w/ alignment (M) achieving up to 81.33% and 82.52% on Detoxify and Perspective API, respectively. These findings highlight that even commercial VLMs are vulnerable to context-driven multimodal attacks, especially when guided by semantically adversarial images. They also underscore the importance of evaluating alignment in continuation-based and visually conditioned scenarios—not just instruction-following tasks.
Table 4 reports the transfer results on LLaMA-3.2-11B-Vision-Instruct, a model with more advanced alignment safeguards. Here, the Text-only Attack yields substantially lower toxicity rates—15.02% under Detoxify and 18.68% under Perspective API—reflecting a high refusal rate for toxic continuations. However, RTD is still able to bypass these safeguards. RTD w/ alignment (L) achieves 41.85% (Detoxify) and 45.24% (Perspective), with RTD (M) performing similarly. These results demonstrate RTD’s ability to bypass even stronger alignment mechanisms, transferring its attack effectiveness across models. Together, they highlight the broader risk of multimodal toxicity attacks against both commercial and open-source VLMs, and underscore the need for more robust, context-aware alignment strategies in future VLM design.
Figure 3. Comparison of images generated by the general-purpose Stable Diffusion (left) and the Red Team Diffuser (right). Given the same image prompts, RTD produces images that subtly elicit discomfort or unease, making them more likely to trigger harmful continuations in VLMs.
Table 5. CLIPScore comparison across different diffusers.
Stable Diffusion RTD (w/o) RTD (L) RTD (M) RTD (H)
CLIPScore 44.47 41.87 42.86 44.02 44.80
5.3. Visualizations
Figure 3 presents a qualitative comparison between images generated by the general-purpose Stable Diffusion model (left) and our Red Team Diffuser (RTD) (right). For RTD-generated images, the alignment reward weights $\lambda$ progressively increase from left to right, starting with a weight of 0 in the leftmost column and followed by 0.05, 0.1, and 0.2. Although all models receive the same image prompts, the resulting images diverge markedly in tone and visual implications. While the general-purpose Stable Diffusion consistently produces neutral and benign visuals. In contrast, RTD generates images that subtly evoke discomfort or toxicity-relevant cues. For instance, in the second row, RTD produces cartoonish and exaggerated caricatures that distort the features of an overweight person eating a sandwich, in contrast to the more neutral depiction by Stable Diffusion. In the third row, where the prompt involves generating an image of a woman, RTD creates visuals that are noticeably more sexually suggestive and explicit. Similarly, in the fourth row, RTD introduces additional elements, such as garbage and ominous shadows, into the school hallway scene generated by Stable Diffusion, thereby amplifying the potential for toxic continuations. The observed differences in visual elements and stylistic features underscore RTD’s ability to adapt visual semantics in ways that exploit the vulnerabilities of target VLMs. The comparison highlights how reinforced optimization—not merely prompt selection—enables subtle but impactful visual manipulation aligned with toxic continuation goals.
To complement these qualitative observations, we report CLIPScore (Hessel et al., 2021) as a quantitative measure of image-text semantic alignment. As shown in Table 5, all RTD variants achieve CLIPScores comparable to or higher than Stable Diffusion, indicating that the adversarial images remain semantically faithful to their prompts despite conveying more toxic or discomforting signals. Notably, RTD with the highest alignment reward ($\lambda=0.2$) attains a CLIPScore of 44.80, surpassing the baseline (44.47) and demonstrating that our stealth-aware reward mechanism enhances both visual regularity and the adversarial effectiveness of the generated images.
6. Limitations
While the Red Team Diffuser (RTD) framework shows strong performance and cross-model transferability, it has several limitations: 1. Task specificity. RTD is designed for toxic continuation tasks, and its effectiveness on other multimodal attacks, such as instruction-level jailbreaks, is yet to be explored. 2. Computational cost. The RL-based fine-tuning requires significant GPU resources, which may be impractical for resource-constrained users.
7. Conclusion
We present Red Team Diffuser (RTD), a novel framework for uncovering critical vulnerabilities in Vision-Language Models (VLMs) through semantically adversarial image generation. We formalize cross-modal toxicity amplification as a new attack surface, where visually coherent yet maliciously tuned images exacerbate harmful text continuations—an aspect largely overlooked by current alignment efforts. RTD combines greedy prompt search with reinforcement fine-tuning to generate images that reliably trigger toxic outputs. Experimental results validate its effectiveness: RTD increases LLaVA’s toxicity rate by 10.69% on the original attack set and 8.91% on a hold-out set, while demonstrating strong cross-model transferability, achieving toxicity gains of 5.1% on Gemini and 26.83% on LLaMA-3.2-11B-Vision-Instruct. Qualitative analysis further illustrates how RTD manipulates visual cues to subtly induce toxic continuations. These findings expose substantial blind spots in existing VLM alignment strategies and underscore the need for more robust, multimodal-aware safety mechanisms. In future work, we aim to expand RTD to broader adversarial objectives and model families, offering deeper insights into VLM vulnerabilities and guiding the development of stronger alignment and defense strategies.